# MemRL: Self-Evolving Agents via Runtime Reinforcement Learning on Episodic Memory

### v1

@@ Abstract
The hallmark of human intelligence is the ability to master new skills through Constructive Episodic Simulation—retrieving past experiences to synthesize solutions for novel tasks. While Large Language Models possess strong reasoning capabilities, they struggle to emulate this self-evolution: fine-tuning is computationally expensive and prone to catastrophic forgetting, while existing memory-based methods rely on passive semantic matching that often retrieves noise. To address these challenges, we propose MemRL, a framework that enables agents to self-evolve via non-parametric reinforcement learning on episodic memory. MemRL explicitly separates the stable reasoning of a frozen LLM from the plastic, evolving memory. Unlike traditional methods, MemRL employs a Two-Phase Retrieval mechanism that filters candidates by semantic relevance and then selects them based on learned Q-values (utility). These utilities are continuously refined via environmental feedback in an trial-and-error manner, allowing the agent to distinguish high-value strategies from similar noise. Extensive experiments on HLE, BigCodeBench, ALFWorld, and Lifelong Agent Bench demonstrate that MemRL significantly outperforms state-of-the-art baselines. Our analysis experiments confirm that MemRL effectively reconciles the stability-plasticity dilemma, enabling continuous runtime improvement without weight updates.
Figure 1: Benchmark Runtime Learning performance of MemRL. We compare MemRL against state-of-the-art memory baselines (MemP) and standard retrieval methods (RAG). MemRL consistently outperforms various baselines, demonstrating the efficacy of runtime utility-driven updates.
@@ 1 Introduction
The hallmark of human intelligence lies in the delicate balance between the stability of cognitive reasoning and the plasticity of episodic memory (grossberg2013adaptive; mcclelland1995there; kumaran2016learning), a mechanism known as Constructive Episodic Simulation that allows adaptation without rewiring neural circuitry (schacter2007constructive; hassabis2007deconstructing; schacter2012future; gick1980analogical). While Large Language Models (LLMs) demonstrate impressive reasoning capabilities, existing paradigms struggle to emulate this dynamic, decoupled self-evolution (wei2022chain; yao2022react; schick2023toolformer; wang2023voyager). On one hand, fine-tuning approaches attempt to internalize experience by modifying model weights (ouyang2022training; stiennon2020learning; rafailov2023direct; ethayarajh2024kto), but often suffer from catastrophic forgetting and high computational costs (kirkpatrick2017overcoming; li2024revisiting; wu2024continuallearninglargelanguage). On the other hand, Retrieval-Augmented Generation (RAG) (lewis2020retrieval) offers a non-parametric alternative but remains fundamentally passive; it retrieves information based solely on semantic similarity without evaluating its actual utility (karpukhin2020dense; gao2023retrieval). Lacking a mechanism to distinguish high-value past strategies from similar noise, current RAG agents struggle to effectively learn from runtime feedback to optimize their performance over time.
@@
This limitation underscores a critical research question: How can we enable an agent to continuously improve its performance after deployment, without compromising the stability of its pre-trained backbone? Our objective is to achieve an agent that evolves with continued usage and rapidly adapts to new tasks after deployment, referred to as Runtime Continuous Learning (javed2023online; silver2025era_of_experience; parisi2019continual; wu2024continuallearninglargelanguage), all while keeping the backbone model frozen to prevent catastrophic forgetting (finn2017model; wei2025evo). To address this challenge, inspired by the human cognitive mechanism of constructive simulation, we propose MemRL, a framework that facilitates self-evolving agents by explicitly decoupling the model’s stable cognitive reasoning from dynamic episodic memory. Figure 2 illustrates the conceptual framework of our proposed MemRL. Drawing on value-iteration approaches in Reinforcement Learning (RL) to estimate expected experience utilities (sutton2018reinforcement), we formalize the interaction between the frozen LLM and external memory as a Markov Decision Process (MDP) (puterman2014markov). Unlike traditional methods that optimize the backbone model, MemRL optimizes the policy of memory usage to maximize expected utility.
@@
MemRL organizes memory into a structured Intent-Experience-Utility triplet. This structure transforms retrieval from a passive semantic match task into an active decision-making process: Value-Aware Retrieval selects experiences based on their learned Q-values, reflecting expected utility, rather than semantic similarity alone (watkins1992q); Utility-Driven Update refines these Q-values via environmental feedback and Bellman backup (bellman1966dynamic). This closed-loop cycle enables the agent to distinguish high-value strategies from similar noise, effectively learning from both success and failure without the computational cost or catastrophic forgetting risks associated with weight updates. As for experiments, we validate MemRL on four diverse benchmarks, including HLE, BigCodeBench, ALFWorld, and Lifelong Agent Bench. Our results demonstrate consistent superiority over baselines, achieving relative improvement in exploration-heavy environments. Our in-depth analysis reveals a strong correlation between learned utility and task success, further confirming MemRL’s effectiveness.
@@
  • We propose a runtime learning framework based on Model-Memory decoupling and the Intent-Experience-Utility triplet, which reconciles the stability-plasticity dilemma by enabling agents to learn without parameter updates.
  • We introduce MemRL, a non-parametric reinforcement learning algorithm that implements Value-Aware Retrieval and Utility-Driven Memory Curation. This allows agents to self-evolve by optimizing memory utility, establishing a new paradigm for enhancing agent capabilities.
  • We conduct extensive evaluations and provide deep insights into MemRL’s working mechanism. We analyze how it ensures structural integrity in complex tasks and theoretically substantiate its stability via Bellman contraction, exploring how utility-driven updates minimize catastrophic forgetting while maximizing positive transfer.
@@ 2 Related Works
2.1 Continuous Learning
Continual learning addresses the stability-plasticity dilemma, aiming to acquire new knowledge sequentially without suffering from catastrophic forgetting. Classical approaches—such as regularization, distillation, and experience replay—mitigate forgetting by constraining parameter updates or preserving past data distributions (kirkpatrick2017overcoming; li2017learning; lopez2017gradient). However, these parametric methods are computationally expensive for LLMs and risk destabilizing the pre-trained backbone through frequent online updates. Recent surveys on continual learning for LLMs further systematize these difficulties and emphasize the importance of external mechanisms and non-parametric pathways (wu2024continuallearninglargelanguage). Therefore, from a continual learning perspective, if we aim for agents to improve with use while preserving the stability of the backbone, a more practical direction is to shift plasticity from the parameter space to external structures and controlled experience-update channels.
2.2 Reinforcement Learning
@@
Reinforcement learning has been widely adopted for LLMs enhancement. A representative paradigm is to construct reward signals from human feedback and optimize the model policy accordingly to align with human preference (stiennon2020learning; ouyang2022training). Other recent approaches leverage rule-based verifiers to improve LLMs’ reasoning capabilities (guo2025deepseek; yu2025dapo). In parallel, agent-oriented research explores how interaction signals can improve tool use and action decision-making, and investigates mechanisms by which language models execute composite actions in environments (schick2023toolformer). Despite the demonstrated effectiveness of reward-driven optimization, these methods generally place learning in the model parameters or additional parametric modules, and thus do not fundamentally avoid the cost of online updates or the risk of forgetting. In contrast, our method frames memory usage as a learnable decision problem and applies non-parametric reinforcement learning on memory to bypass the risk.
2.3 Agentic Memory
@@
To bypass the costs of fine-tuning, external memory systems have evolved from a static RAG paradigm to dynamic, governable memory structures (lewis2020retrieval; karpukhin2020dense). Early agentic memory introduced reflection mechanisms and hierarchical management to handle long context experiences (shinn2023reflexion; packer2024memgptllmsoperatingsystems). More recent frameworks have systematized the memory lifecycle, focusing on unified storage and structured indexing for complex tasks (li2025memos; xu2025mem; huang2025licomemory; ye2025task). Furthermore, adaptive approaches now explore improving retrieval via feedback-driven updates or automated augmentation (salama2025meminsight; zhang2025learn; li2025retrieval; zhou2025memento). However, except for training additional learnable modules, most existing methods still rely predominantly on semantic similarity or heuristic rules, lacking a rigorous metric to evaluate the actual utility of a memory in maximizing returns. Inspired by cognitive theories of memory reconsolidation (schacter2007constructive; gick1980analogical; nader2000fear), MemRL bridges this gap by formulating retrieval as a value-based decision process, learning robust utility estimates (Q-values) directly from environmental rewards to distinguish high-value experiences from noise.
@@ 3 Problem Formulation
In this section, we formally define the problem of memory-augmented generation and establish the theoretical link between agent policy and memory retrieval. We adopt the formulation of Memory-Based Markov Decision Process (M-MDP) (zhou2025memento) and address it with our non-parametric reinforcement learning framework. Figure 3 provides an illustrative example of this memory-augmented decision process, showing how retrieval outcomes and memory evolution unfold over multiple time steps.
Figure 3: An illustrative example of memory-augmented decision making under a Markov Decision Process. At time step $t$, the agent starts with an initial memory set $\mathcal{M}_{t}$. At time step t+1t{+}1, an intent (Intent A) retrieves relevant past experiences, but initially leads to a failed generation. In contrast, another intent (Intent B) succeeds and its associated experience is added to memory. At time step t+2t{+}2, Intent A retrieves the newly stored successful experience from Intent B, resulting in a successful outcome. This example shows how memory retrieval enables knowledge reuse across intents, implicitly supporting transfer across tasks through shared experiences.
@@ 3.1 Memory-Augmented Agent Policy
To enable the agent to self-evolve, we inherit the M-MDP as the problem formulation (zhou2025memento). The M-MDP is formally defined as a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{M}\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ the transition dynamics, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1)$ the discount factor, and $\mathcal{M}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{*}$ the evolving memory space containing past experiences (zhou2025memento). In this setting, the policy $\pi$ not only generates tokens but first selects a memory context $m$ based on a retrieval distribution $\mu(\cdot|s,\mathcal{M})$, allowing the agent to leverage historical data for better performance in downstream tasks.
We consider a generalist agent interacting with an environment or user over discrete time steps. At each step $t$, the agent receives a state $s_{t}$, e.g., a user query or task description, and has access to an external memory bank $\mathcal{M}$. The agent’s objective is to generate a response $y_{t}$ that maximizes a reward signal. Following this formulation, the behavior of a memory-augmented agent can be decomposed into two distinct phases: Retrieve and Generation. The joint policy $\pi(y_{t}|s_{t},\mathcal{M}_{t})$ is defined as the marginal probability over all possible retrieved memory items (zhou2025memento):
$\pi(y_{t}|s_{t},\mathcal{M}_{t})=\sum_{m\in\mathcal{M}_{t}}\underbrace{\mu(m|s_{t},\mathcal{M}_{t})}_{\text{Retrieval Policy}}\cdot\underbrace{p_{\text{LLM}}(y_{t}|s_{t},m)}_{\text{Inference Policy}}$ (1)
where:
  • $\mu(m|s_{t},\mathcal{M}_{t})$ represents the Retrieval Policy, which assigns a probability to selecting a specific memory context $m$ composed of past intents and experiences. from the memory bank $\mathcal{M}_{t}$ given the current state $s_{t}$.
  • $p_{\text{LLM}}(y_{t}|s_{t},m)$ represents the Inference Policy, typically parameterized by a frozen LLM. It models the likelihood of generating output $y_{t}$ conditioned on both the query $s_{t}$ and the retrieved context $m$.
In previous RAG or memory-based agentic paradigms, the retrieval policy $\mu$ is usually determined by a fixed vector similarity metric, e.g., cosine similarity of embeddings. While effective for semantic matching, such policies fail to account for the utility of a memory, i.e., whether retrieving $m$ actually leads to a successful outcome $y_{t}$.
@@ 3.2 Non-Parametric Reinforcement Learning
To overcome the limitations of static similarity-based retrieval, we operationalize the M-MDP framework by formulating memory retrieval as a value-based decision-making process. Unlike parametric approaches that optimize $\pi_{\text{LLM}}$ via weight updates, we aim to optimize the retrieval policy $\mu(m|s,\mathcal{M})$ directly within the memory space. We map the abstract M-MDP components to a specific Intent-Experience-Utility structure:
From Semantic Matching to Decision Making. We instantiate the state $s$ as the User Intent, encapsulated by the embedding of the current query (lewis2020retrieval). Consequently, the action space $\mathcal{A}_{t}$ becomes dynamic and discrete, corresponding to the selection of a specific $m$ from the current memory bank $\mathcal{M}_{t}$. Under this formulation, retrieving from memory is no longer a passive matching task but an active action $a_{t}=m$ taken to augment the generator (zhou2025memento).
Defining Utility via Q-Values. The core of our framework is the shift from estimating semantic relevance to estimating functional utility. We define the state-action value function $Q(s,m)$ as the expected utility of applying the retrieved context $m$ for intents similar to $s$. The objective of MemRL is to learn an optimal retrieval policy $\mu^{*}$ that selects context maximizing this expected utility:
$\mu^{*}(m|s,\mathcal{M})=\arg\max_{m\in\mathcal{M}}Q(s,m)$ (2)
This $Q$-value serves as a critic, distinguishing high-value strategies from irrelevant noise that may share high semantic similarity.
Non-Parametric Learning. Since the action space of the retrieval policy $\mu$ in our setting is decoupled from the LLMs’ generation space, we can perform learning without modifying the LLMs’ weights. Upon receiving an environmental feedback $r$, e.g., execution success, we can update the Q-value of the retrieved memory context using a Temporal-Difference (TD) error (sutton1988learning):
$Q(s,m)\leftarrow Q(s,m)+\alpha[r+\gamma\max Q(s^{\prime},m^{\prime})-Q(s,m)]$ (3)
where $\alpha$ is the learning rate. This Bellman-style backup allows the utility estimates to converge to the true expected returns over time (bellman1966dynamic). By maintaining and updating these Q-values explicitly within the memory structure, MemRL provides a non-parametric learning manner with a theoretical guarantee that enables the agent to self-evolve its capabilities through interaction.

### v2

@@ Abstract
The hallmark of human intelligence is the self-evolving ability to master new skills by learning from past experiences. However, current AI agents struggle to emulate this self-evolution: fine-tuning is computationally expensive and prone to catastrophic forgetting, while existing memory-based methods rely on passive semantic matching that often retrieves noise. To address these challenges, we propose MemRL, a non-parametric approach that evolves via reinforcement learning on episodic memory. By decoupling stable reasoning from plastic memory, MemRL employs a Two-Phase Retrieval mechanism to filter noise and identify high-utility strategies through environmental feedback. Extensive experiments on HLE, BigCodeBench, ALFWorld, and Lifelong Agent Bench demonstrate that MemRL significantly outperforms state-of-the-art baselines, confirming that MemRL effectively reconciles the stability-plasticity dilemma, enabling continuous runtime improvement without weight updates. Code is available at https://github.com/MemTensor/MemRL.
Machine Learning, ICML
@@ 1 Introduction
Human intelligence balances cognitive stability and episodic plasticity (Grossberg, 2013; McClelland et al., 1995; Kumaran et al., 2016) via Constructive Episodic Simulation, enabling adaptation without rewiring neural circuitry (Schacter and Addis, 2007; Hassabis and Maguire, 2007; Schacter et al., 2012; Gick and Holyoak, 1980). Despite their reasoning capabilities, current AI agents struggle to emulate this decoupled self-evolution (Wei et al., 2022; Yao et al., 2023; Schick et al., 2023; Wang et al., 2023). Specifically, fine-tuning internalizes experience by modifying weights (Ouyang et al., 2022; Stiennon et al., 2020; Rafailov et al., 2023; Ethayarajh et al., 2024) but suffers from computational costs and catastrophic forgetting (Kirkpatrick et al., 2017; Li et al., 2024; Wu et al., 2024). Conversely, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) provides a non-parametric alternative but remains passive, retrieving by semantic similarity rather than utility (Karpukhin et al., 2020; Gao et al., 2024); this prevents agents from effectively leveraging runtime feedback to distinguish high-value strategies from noise.
@@
This limitation underscores a critical research question: How can we enable an agent to continuously improve its performance after deployment, without compromising the stability of its pre-trained backbone? Our objective is to achieve an agent that evolves with continued usage and rapidly adapts to new tasks after deployment, referred to as Runtime Continuous Learning (Javed et al., 2023; Silver and Sutton, 2025; Parisi et al., 2019; Wu et al., 2024), all while keeping the backbone model frozen to prevent catastrophic forgetting (Finn et al., 2017; Wei et al., 2025). To address this challenge, inspired by the human cognitive mechanism of constructive simulation, we propose MemRL, an approach that facilitates self-evolving agents by explicitly decoupling the model’s stable cognitive reasoning from dynamic episodic memory. Figure 1 illustrates the conceptual framework of our proposed MemRL. Drawing on tries-and-errors manner in Reinforcement Learning (RL) to estimate expected experience utilities (Sutton and Barto, 2018), we formalize the interaction between the frozen LLM and external memory as a Markov Decision Process (MDP) (Puterman, 2014). Unlike traditional methods that optimize the backbone model, MemRL optimizes the policy of memory usage without tuning model weights.
@@
MemRL organizes memory into a structured Intent-Experience-Utility triplet. This structure transforms retrieval from a passive semantic match task into an active decision-making process: Two-Phase Retrieval selects experiences based on their learned Q-values, reflecting expected utility, rather than semantic similarity alone (Watkins and Dayan, 1992); Utility-Driven Update refines these Q-values through environmental feedback, applying Monte Carlo style updates (Metropolis and Ulam, 1949). This closed-loop cycle enables the agent to distinguish high-value memories from similar noise, effectively learning from both success and failure without high computational cost or catastrophic forgetting risks associated with weight updates. As for experiments, we validate MemRL on four diverse benchmarks, including HLE, BigCodeBench, ALFWorld, and Lifelong Agent Bench. Our results demonstrate consistent superiority over baselines, achieving relative improvement in exploration-heavy environments. Our in-depth analysis reveals a strong correlation between learned utility and task success, further confirming MemRL’s effectiveness.
@@
  • We propose a runtime learning framework using Model-Memory decoupling and Intent-Experience-Utility triplet to reconcile the stability-plasticity dilemma, enabling tuning-free agent learning.
  • We introduce MemRL, a non-parametric approach enabling agent self-evolution via Two-Phase Retrieval and Utility-Driven Update.
  • We conduct extensive evaluations and provide a rigorous analysis for MemRL’s stability, showing how it ensures task integrity and minimizes forgetting.
@@ 2 Related Works
Runtime Learning
Runtime Learning focuses on the post-deployment improvement of agents through interaction streams rather than offline data, marking a shift toward the “era of experience” (Silver and Sutton, 2025). Unlike Continual Learning (Parisi et al., 2019; Wu et al., 2024) or Test-Time Adaptation (Sun et al., 2020; Wang et al., 2020; Liang et al., 2025), which typically update parameters to handle forgetting or distribution shifts, our setting constrains the backbone to remain frozen to ensure stability and efficiency. While recent memory-augmented agents (Zheng et al., 2025; Wei et al., 2025; Zhou et al., 2025b) emphasize memory organization, the selection problem—identifying which experiences to reuse under feedback—remains a critical challenge. Drawing from value-aware episodic control (Tulving and others, 1972; McClelland et al., 1995), we frame runtime learning as identifying valuable episodes. By using interaction feedback to assign utility, our approach guides retrieval and reuse without weight modification, thereby ensuring sustained improvement (Pritzel et al., 2017).
Reinforcement Learning
@@
Reinforcement learning has been widely adopted for LLMs enhancement. A representative paradigm is to construct reward signals from human feedback and optimize the model policy accordingly to align with human preference (Stiennon et al., 2020; Ouyang et al., 2022). Other recent approaches leverage rule-based verifiers to improve LLMs’ reasoning capabilities (Guo et al., 2025; Yu et al., 2025). In parallel, agent-oriented research explores how interaction signals can improve tool use and action decision-making, and investigates mechanisms by which language models execute composite actions in environments (Schick et al., 2023). Despite the demonstrated effectiveness of reward-driven optimization, these methods generally place learning in the model parameters or additional parametric modules, and thus do not avoid the cost of online updates or the risk of forgetting. In contrast, our method frames memory usage as a learnable decision problem and applies non-parametric reinforcement learning on memory to bypass the risk.
Agentic Memory
Figure 2: An illustrative example of memory-augmented decision making under a Markov Decision Process. At time step $t$, the agent starts with an initial memory set $\mathcal{M}_{t}$. At time step t+1t{+}1, an intent (Intent A) retrieves relevant past experiences, but initially leads to a failed generation. In contrast, another intent (Intent B) succeeds, and its associated experience is added to memory. At time step t+2t{+}2, Intent A retrieves the newly stored successful experience from Intent B, resulting in a successful outcome. This example shows how memory retrieval enables knowledge reuse across intents, implicitly supporting self-evolution through shared experiences.
@@
To avoid the costs of fine-tuning, external memory systems have evolved from a static RAG paradigm to dynamic, governable memory structures (Lewis et al., 2020; Karpukhin et al., 2020). Early agentic memory introduced reflection mechanisms and hierarchical management to handle long context experiences (Shinn et al., 2023; Packer et al., 2024). More recent frameworks have systematized the memory lifecycle, focusing on unified storage and structured indexing for complex tasks (Li et al., 2025b; Xu et al., 2025; Huang et al., 2025; Ye, 2025). Furthermore, adaptive approaches now explore improving retrieval via feedback-driven updates or automated augmentation (Salama et al., 2025; Zhang et al., 2025; Li et al., 2025a; Zhou et al., 2025a). However, except for training additional learnable modules, most existing methods still rely predominantly on semantic similarity or heuristic rules, lacking a rigorous metric to evaluate the actual utility of a memory in maximizing returns. Inspired by cognitive theories of memory reconsolidation (Schacter and Addis, 2007; Gick and Holyoak, 1980; Nader et al., 2000), MemRL bridges this gap by formulating retrieval as a value-based decision process, learning robust utility estimates (Q-values) from environmental rewards to distinguish high-value experiences.
@@ 3 Problem Formulation
In this section, we formally define the problem of memory-augmented generation and establish the theoretical link between agent policy and memory retrieval. We adopt the formulation of Memory-Based Markov Decision Process (M-MDP) (Zhou et al., 2025a), and apply our non-parametric reinforcement learning approach to it. Figure 2 provides an illustrative example of this memory-augmented decision process, showing how retrieval outcomes and memory evolution unfold over multiple time steps.
@@ 3.1 Memory-Augmented Agent Policy
To enable the agent to self-evolution, we adopt the M-MDP framework (Zhou et al., 2025a), defined by the tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R},\gamma,\mathcal{M})$. Here, $\mathcal{S}$ and $\mathcal{A}$ represent state and action spaces, $P$ is the transition dynamics, $\mathcal{R}$ is the reward function of state and action, $\gamma\in[0,1)$ denotes the discount factor, and $\mathcal{M}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{*}$ constitutes the evolving memory bank of past experiences (Zhou et al., 2025a). At each step $t$, the agent receives state $s_{t}$ and leverages $\mathcal{M}_{t}$ to generate a response $a_{t}$ maximizing the expected reward. The joint policy $\pi(a_{t}|s_{t},\mathcal{M}_{t})$ is formulated as the marginal probability over all possible retrieved items $m$ (Zhou et al., 2025a):
$\pi(a_{t}|s_{t},\mathcal{M}_{t})=\sum_{m\in\mathcal{M}_{t}}\mu(m|s_{t},\mathcal{M}_{t})p_{LLM}(a_{t}|s_{t},m).$ (1)
where $\mu(m|s_{t},\mathcal{M}_{t})$ is the Retrieval Policy for selecting memory contexts, and $p_{LLM}(a_{t}|s_{t},m)$ is the Inference Policy parameterized by a frozen LLM. This approach transforms retrieval from a passive match into an active decision process, effectively accounting for the functional utility of $m$ in generating successful outcomes $a_{t}$.
In previous RAG or memory-based agentic paradigms, the retrieval policy $\mu$ is usually determined by a fixed vector similarity metric, e.g., cosine similarity of embeddings. While effective for semantic matching, such policies fail to account for the utility of a memory, i.e., whether retrieving $m$ actually leads to a successful outcome $a_{t}$.
Figure 3: Overview of MemRL. The end-to-end learning loop(Top): given a query $\mathbf{s}$, the agent retrieves context $\mathbf{m}_{ctx}$ from memory $\mathbf{M}$, generates output $\mathbf{y}$, and updates memory value $Q$ based on reward $R$. Two-Phase Retrieval(Bottom Left): Candidates are recalled via similarity, then re-ranked using learned Q-values. Utility Update(Bottom Right): Values ($Q$) are updated using environmental rewards to distinguish functional utility from semantic similarity.
@@ 3.2 Non-Parametric Reinforcement Learning
To overcome static similarity limitations, we operationalize the M-MDP framework by formulating memory retrieval as a value-based decision-making process (Zhou et al., 2025a). Unlike parametric methods optimizing $\pi_{LLM}$ via weight updates, we optimize the retrieval policy $\mu(m|s,\mathcal{M})$ directly within the memory space by mapping M-MDP components to a structured Intent-Experience-Utility triplet:
From Semantic Matching to Decision Making. We instantiate state $s$ as the User Intent, encapsulated by the current query embedding (Lewis et al., 2020). Consequently, the action space $\mathcal{A}_{t}$ becomes dynamic and discrete, corresponding to selecting a specific $m$ from the memory bank $\mathcal{M}_{t}$ (Zhou et al., 2025a). In this formulation, retrieval is not a passive matching task but a strategic decision step taken to augment the generator’s action $a$ (Zhou et al., 2025a).
Defining Utility via Q-Values. While the agent’s executable action $a$ is generated by the policy $\pi$ (as formalized in Sec. 3.1), the quality of this generation is strictly conditioned on the retrieved context. Therefore, we adapt the traditional value function $Q(s,a)$ to the retrieval phase, defining $Q(s,m)$ as the expected utility of the subsequent action $a$ augmented by memory $m$. MemRL learns an optimal retrieval policy $\mu^{*}$ that maximizes this utility:
$\mu^{*}(m|s,\mathcal{M})=\arg\max_{m\in\mathcal{M}}Q(s,m).$ (2)
In this view, the Q-value acts as a critic for the retrieval mechanism, distinguishing memories that strategically aid the generator from irrelevant noise that merely shares high semantic similarity.
Non-Parametric Learning. Since the retrieval action space is decoupled from LLM generation, we perform learning without modifying model weights. Upon receiving environmental feedback $r$, we update the Q-value via a Temporal-Difference (TD) error (Sutton, 1988):
$Q(s,m)\leftarrow Q(s,m)+\alpha[r+\gamma\max Q(s^{\prime},m^{\prime})-Q(s,m)],$ (3)
or Monte Carlo style rule (Metropolis and Ulam, 1949):
$Q_{\text{new}}\leftarrow Q_{\text{old}}+\alpha\big(r-Q_{\text{old}}\big),$ (4)
where $\alpha$ is the learning rate. Equation 4 performs as a naturally simplified version of Equation 3 by setting $s^{\prime}$ as a terminal state to balance complexity and performance, sharing a similar one-step MDP formulation with Guo et al. (2025). These manners allow utility estimates to converge to true expected returns over time (Bellman, 1966). By explicitly updating Q-values within the memory structure, MemRL provides a non-parametric learning manner with a theoretical guarantee, enabling agents to self-evolve through interaction.